OpenStudy (anonymous):

f(x)=x^(x/2) what is f'(x)?

OpenStudy (anonymous):

\[f(x)=x^{x/2}\]

OpenStudy (anonymous):

Oh this is a nice one. Do you know logarithmic differentiation?

OpenStudy (anonymous):

say y=x^(x/2) ln y =x/2 lnx d(ln y)=d(x ln x)/2 y'/y=1/2 * (ln x+x*1/x) y'/y=1/2*(ln x+1) y'=y/2 * (ln x+1) y'=x^(x/2)*(ln x+1)

OpenStudy (anonymous):

Sorry, I shouldn't post the answer right away. But I hope this helps :)

OpenStudy (anonymous):

And i missed our 1/2 at the final answer.

OpenStudy (anonymous):

@imron07 you forgot the 1/2 in the final result.

OpenStudy (anonymous):

hehe, thanks @gezimbasha

OpenStudy (anonymous):

thanks. but how did d(ln y)= y'/y ?

OpenStudy (anonymous):

Actually, d(ln x)/dx =1/x For u(x), use chain rule: d(ln u)/dx=1/u * du/dx =u'/u

OpenStudy (anonymous):

oh. ok. thanks. :)

OpenStudy (anonymous):

Okay :)